OpenStudy (anonymous):

What is the value of the expression 14 + 6 • 2 – 10 ÷ 5? A. 0 B. 38 C. 6 D. 24

OpenStudy (trojanpoem):

0

OpenStudy (trojanpoem):

20/2 = 10 10 - 10 = 0 /5 = 0

OpenStudy (anonymous):

Thank you